OpenStudy (anonymous):

P ( -1/3 ,1/5) . Find Cos(theta)

OpenStudy (anonymous):

\[P ( \frac{ -1 }{ 3 } ,\frac{ 1 }{ 5 } ) . Find \cos \]

OpenStudy (zepp):

What does this P thing means?

OpenStudy (anonymous):

Point in circular

OpenStudy (anonymous):

i mean point on circle (x,y )

OpenStudy (zepp):

Oh, I see

OpenStudy (zepp):

 Like this?

OpenStudy (anonymous):

idk :|

OpenStudy (anonymous):

and how to find theta

OpenStudy (zepp):

 ;D Go ahead and solve! There coordinates are a big hint of the length of these triangles sides.